You have 188 feet of fencing to enclose a rectangular region. What is the maximum area?
Answer by rapaljer(4671)

The maximum area enclosed in a rectangular region is a square--GUARANTEED!

So, P= 4s
4s= 188
s= 47 feet

A= s^2
A= 47^2 = 2209 square feet

R^2 at SCC